7-2A - CVP Analysis, What if? Analysis Hewins Incs projected contribution-format income statement for the upcoming year is
Answer #1
Total Per unit (total/no. of units)
No. of units sold 10000
Sales /SP $       2,000,000.00 $                       200.00
Variable expenses / VC $       1,400,000.00 $                       140.00
Contribution margin $           600,000.00 $                         60.00
Fixed expenses $           500,000.00
Net operating income $           100,000.00 $                         10.00
a) Break even point in units =fixed expenses / (SP - VC)
=500,000 / 60
8334 units
b) Break even point in dollars = Fixed costs / [(SP-VC)/SP]
=500,000/[(200-140)/200]
$       1,666,666.67
c) Target profit $300,000 when target profit is given add it to the fixed cost and calculate the Break even point
=(fixed expenses + profit) / (SP - VC)
=(500,000+300,000) / 60
13334 units
d)Margin of safety it shows the total no. of sales above break even
in units Current Sales Units - Breakeven Sale Units
=10,000 - 8334
1666 Units
in dollars Current Sales - Break Even Sale
=2,000,000 - 1,666,667
$           333,333.33